Find the volume of the tetrahedron boulded by the coordinate planes and theplane x/a+y/b+z/c=1


1
Expert's answer
2021-09-14T01:53:56-0400

The volume bounded by coordinates plane and the plan x/a+y/b+z/c =1​ is abc/6 cubic.units.

To find: the volume bounded by coordinates plane and the plan x/a+y/b+z/c =1​

Given,

The equations of the lines situated on the planes y=0 and z=0 are:

y=0 gives,

x/a+z/c = 1 ⇔ z = (1−x/a)c,

And,

z=0 gives,

x/a+y/c=1 ⇔ y=(1−x/a)b.

Area A(x) is given by,

A(x) = 1/2(1−x/a)b(1−x/a)c

= bc/2(1−x/a)^2.

Hence the volume is given by the integration of the area A(x) from x = 0 to x = a





V = bc/2(a − 2/a × a^2/2 + 1/a^2 × a^3/3)


V = abc/6.